# IMU: Influence-guided Machine Unlearning

### v1

@@ 6.3 Derivation of influence function
0 0 $\approx\nabla_{{\bm{\theta}}}R({\bm{\theta}}^{\ast})+\epsilon\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{\theta}}^{\ast})+\left[\nabla_{{\bm{\theta}}}^{2}R({\bm{\theta}}^{\ast})+\epsilon\nabla_{{\bm{\theta}}}^{2}\ell({\bm{x}},{\bm{\theta}}^{\ast})\right]\Delta_{\epsilon}.$ (29)
@@ 7.1 Results on sample-wise unlearning
In Table 7, we evaluate the unlearning performance of IMU method on ResNet-18 for random forgetting on CIFAR-10. The results clearly manifest that our proposed approach achieves robust balance across these metrics with feasible run time, compared with other existing methods. Additionally, we also examine how different influence update frequencies affect sample-wise forgetting performance in Table 6. And we can also develop the conclusion that in this scenario, even infrequent updates are sufficient for effective unlearning.
@@
$\nu$ $\texttt{Acc}_{{\mathcal{D}}_{f}}(\downarrow)$ $\texttt{Acc}_{{\mathcal{D}}_{r}}(\uparrow)$ $\texttt{Acc}_{{\mathcal{D}}_{t}}(\uparrow)$ MIA ()($\uparrow$)( ↑ ) Run time ()($\downarrow$)( ↓ ) 0 97.82$\pm$0.31 97.78$\pm$0.53 91.90$\pm$0.82 0.12$\pm$0.06 055$\pm$0 111 97.47$\pm$0.42 97.55$\pm$0.49 91.82$\pm$0. 0.08$\pm$0.03 224$\pm$1 222 97.17$\pm$0.00 97.17$\pm$0.97 91.50$\pm$1.07 0.11$\pm$0.03 139$\pm$1
@@
Setting Method ${\mathcal{D}}_{r}$ ${\mathcal{D}}_{f}$ $\texttt{Acc}_{{\mathcal{D}}_{f}}(\downarrow)$ $\texttt{ACC}_{{\mathcal{D}}_{r}}(\uparrow)$ $\texttt{ACC}_{{\mathcal{D}}_{t}}(\uparrow)$ MIA ()($\uparrow$)( ↑ ) $W_{\texttt{dist}}(\downarrow)$ Run time (s) ()($\downarrow$)( ↓ ) CIFAR-10 Original 099.47$\pm$0.00 100.00$\pm$0.00 094.61$\pm$0.00 00.00$\pm$0.00 Retrain 094.33$\pm$0.00 099.93$\pm$0.00 094.42$\pm$0.00 00.13$\pm$0.00 00.00$\pm$0.00 IU 099.06$\pm$0.11 099.03$\pm$0.03 092.98$\pm$0.06 00.02$\pm$0.01 00.06$\pm$0.04 054$\pm$1 GA 099.04$\pm$0.47 098.75$\pm$0.27 092.88$\pm$0.18 00.02$\pm$0.00 00.07$\pm$0.04 031$\pm$1 RL 098.90$\pm$0.55 098.90$\pm$0.51 093.49$\pm$0.55 00.11$\pm$0.00 00.03$\pm$0.01 024$\pm$0 SSD 099.04$\pm$0.44 099.01$\pm$0.36 093.34$\pm$0.54 00.03$\pm$0.00 00.04$\pm$0.01 055$\pm$3 SCAR 098.78$\pm$0.25 098.94$\pm$0.14 093.17$\pm$0.24 00.13$\pm$0.02 00.04$\pm$0.01 703$\pm$2 NPO 098.74$\pm$0.16 098.83$\pm$0.15 092.98$\pm$0.22 00.03$\pm$0.02 00.05$\pm$0.03 038$\pm$1 IMU (ours) 098.64$\pm$0.11 099.06$\pm$0.20 093.62$\pm$0.21 00.03$\pm$0.00 00.08$\pm$0.01 035$\pm$0 CIFAR-100 Original 097.69$\pm$0.00 097.48$\pm$0.00 076.25$\pm$0.00 00.06$\pm$0.00 Retrain 075.71$\pm$0.00 099.98$\pm$0.00 074.24$\pm$0.00 00.50$\pm$0.00 00.00$\pm$0.00 IU 094.58$\pm$1.62 094.71$\pm$1.76 070.11$\pm$0.77 00.12$\pm$0.01 00.67$\pm$0.21 057$\pm$1 GA 094.42$\pm$0.93 093.36$\pm$0.57 070.37$\pm$0.15 00.11$\pm$0.01 01.07$\pm$0.51 032$\pm$1 RL 093.82$\pm$1.46 094.14$\pm$1.51 066.58$\pm$1.05 00.06$\pm$0.03 01.07$\pm$0.18 031$\pm$1 SSD 093.97$\pm$0.73 094.22$\pm$0.80 070.30$\pm$0.23 00.13$\pm$0.01 00.31$\pm$0.14 050$\pm$0 SCAR 093.82$\pm$1.04 095.18$\pm$1.03 070.70$\pm$0.60 00.14$\pm$0.01 00.81$\pm$0.11 726$\pm$4 NPO 094.43$\pm$0.70 094.31$\pm$0.53 070.86$\pm$0.60 00.12$\pm$0.00 01.02$\pm$0.25 038$\pm$2 IMU (ours) 093.55$\pm$0.23 095.21$\pm$0.09 072.15$\pm$0.88 00.14$\pm$0.01 00.74$\pm$0.31 062$\pm$2
@@ 7.2 Results on LLM unlearning
In Table 8, we generalize our MU method on large language model and perform forgetting forget05 task on the TOFU benchmark. Compared with other LLM unlearning baselines, IMU (ours) strikes a balance on forget quality and model utility. From Figure , we can clearly see the catastrophic collapse issue in the GA method and better performance of the NPO. Although the model structure of Llama-3.2-3B-Instruct is more complex, which is not conducive to the accurate estimation of influence function, our method has achieved an almost equivalent forgetting effect to SimNPO method.
Table 8: Results of LLM unlearning.
@@
GA 0.00 0.03 0.03 0.00 480
NPO 0.28 0.05 0.06 0.21 800
SimNPO 0.35 0.05 0.17 0.33 -120
IMU (Ours) 0.33 0.06 0.09 0.29 -830
@@ 7.3 Experiments on SalUn
Under the same dataset and experimental settings, we replicated and evaluated the current SOTA method based on the retained dataset and forgotten dataset SalUn (Fan et al. 2024b). The experimental results are shown in the Table 9. On the CIFAR-10 dataset, the IMU method we proposed performs similarly to SalUn, while on the CIFAR-100 dataset, SalUn still achieves relatively better performance.
@@
Setting $\texttt{Acc}_{{\mathcal{D}}_{f}}(\downarrow)$ $\texttt{Acc}_{{\mathcal{D}}_{r}}(\uparrow)$ $\texttt{Acc}_{{\mathcal{D}}_{t}}(\uparrow)$ MIA ()($\uparrow$)( ↑ ) Run time ()($\downarrow$)( ↓ )
CIFAR-10 single class 00.07$\pm$0.18 99.40$\pm$0.39 87.29$\pm$0.48 1.00$\pm$0.00 151$\pm$2
CIFAR-100 subclass 00.00$\pm$0.00 96.80$\pm$4.05 83.40$\pm$9.17 1.00$\pm$0.00 024$\pm$0
CIFAR-100 subclass 00.00$\pm$0.00 99.61$\pm$0.47 91.35$\pm$2.83 1.00$\pm$0.00 029$\pm$4
CIFAR-10 random 97.35$\pm$0.23 98.98$\pm$0.16 93.03$\pm$0.15 0.12$\pm$0.00 237$\pm$3
CIFAR-100 random 78.27$\pm$1.08 98.85$\pm$0.42 70.03$\pm$0.78 0.69$\pm$0.02 249$\pm$6
@@ 7.4 Changes of influence score over epochs
In Figure 4, we respectively present the changes in the influence scores of 100100100 samples randomly selected from ${\mathcal{D}}_{f}$ over the 555 unlearning epochs in the class-wise and sample-wise forgetting tasks. In the class-wise forgetting task, since the samples are all belong to the same class, their values should all be negative according to the definition of the influence function formula. It is verified in Figure 4, and we can see the changes in the influence score of a specific sample after epoch 222 do not significantly fluctuate. Additionally, compared to the class-wise task, since the samples are randomly selected from different classes in the sample-wise forgetting scenario, the overall influence of each sample on forgetting data is possibly positive or negative. We can also observe that with more forgetting rounds, the number of samples with negative influence values gradually increases, demonstrating the effectiveness of our unlearning method.
@@
Figure 4: Change of sample influences for 555 epochs in CIFAR-10, the left column represents the class-wise forgetting tasks and the right column represents the random-wise forgetting tasks.

### v2

@@ 6.3 Derivation of influence function
$0$ $\approx\nabla_{{\bm{\theta}}}R({\bm{\theta}}^{\ast})+\epsilon\nabla_{{\bm{\theta}}}\ell({\bm{x}},{\bm{\theta}}^{\ast})+\left[\nabla_{{\bm{\theta}}}^{2}R({\bm{\theta}}^{\ast})+\epsilon\nabla_{{\bm{\theta}}}^{2}\ell({\bm{x}},{\bm{\theta}}^{\ast})\right]\Delta_{\epsilon}.$ (29)
@@ 7.1 Results on sample-wise unlearning
In Table 7, we evaluate the unlearning performance of IMU method on ResNet-18 for sample-wise unlearning on CIFAR-10. The results clearly manifest that our proposed approach achieves robust balance across these metrics with feasible run time, compared with other existing methods. Additionally, we also examine how different influence update frequencies affect sample-wise unlearning performance in Table 6. And we can also develop the conclusion that in this scenario, infrequent updates are sufficient for effective unlearning, which validates the robustness of our method.
$\nu$ $\texttt{Acc}_{{\mathcal{D}}_{f}}(\downarrow)$ $\texttt{Acc}_{{\mathcal{D}}_{r}}(\uparrow)$ $\texttt{Acc}_{{\mathcal{D}}_{t}}(\uparrow)$ MIA $(\uparrow)$ Run time $(\downarrow)$
[…]
@@
[…]
@@ 7.2 Results on LLM unlearning
In Table 8 and Figure 4, we generalize our MU method on large language model and perform forgetting forget05 task on the TOFU benchmark. Compared with other LLM unlearning baselines, IMU (ours) strikes a balance on forget quality and model utility. From Figure 4, we can clearly see the catastrophic collapse issue in the GA method and slightly better performance of the NPO. Although the model structure of Llama-3.2-3B-Instruct is more complex than that of image task, which is not conducive to the accurate estimation of influence function, our method has still achieved an almost equivalent forgetting effect to SimNPO method.
Figure 4: Comparison of forget quality and model utility between GA, NPO, SimNPO, and IMU.
In Table 10, We presented $5$ randomly selected question-answer pairs from forget05 of TOFU benchmark and the corresponding output answers for different unlearning methods. Among them, the GA method always chooses not to answer or simply repeats the questions for difficult questions. Even in a simple question Q3, it did not forget the key information “Global Health”. In contrast, although NPO has also produces some meaningless information, in general, it ensured the fluency and logic of the output language. While the SimNPO method performed well in forgetting important information, but it still retained related information similar to the ground truth in some responses. However, our method IMU, despite significantly forgetting all relevant information, still maintained effective output, which greatly enhancing fluency and diversity of the generated content.
In Table 11, the question-answer pairs were selected from retain95. It means that the output of the model should be approximate to the ground truth, which proves the retention performance of the unlearned model. From Table 11, it is clearly observable that the answer provided by GA is always very simple and lacks detailed descriptions. Although NPO and SimNO’s answers are relatively accurate, they sometimes also output information that deviates from the ground truth. In contrast, our method is the closest to the facts, demonstrating the best model utility among these unlearn methods.
@@
GA 0.00 0.03 0.03 0.00 048
NPO 0.28 0.05 0.06 0.21 080
SimNPO 0.35 0.05 0.17 0.33 -12
IMU (Ours) 0.33 0.06 0.09 0.29 -83
Table 8: Results of LLM unlearning on TOFU (Forget05) over 66 epochs.
@@ 7.3 Experiments on SalUn
Under the same dataset and experimental settings, we replicated and evaluated the current SOTA method based on the retain dataset and forget dataset SalUn (Fan et al. 2024b). The experimental results are shown in Table 9. On the CIFAR-10 dataset, the result of IMU method we proposed performs is similarly to SalUn, while on the CIFAR-100 dataset, SalUn still achieves relatively better performance.
Setting $\texttt{Acc}_{{\mathcal{D}}_{f}}(\downarrow)$ $\texttt{Acc}_{{\mathcal{D}}_{r}}(\uparrow)$ $\texttt{Acc}_{{\mathcal{D}}_{t}}(\uparrow)$ MIA $(\uparrow)$ Run time $(\downarrow)$
[…]
@@ 7.4 Changes of influence score over epochs
In Figure 5, we respectively present the changes in the influence scores of $100$ samples randomly selected from ${\mathcal{D}}_{f}$ over the $5$ unlearning epochs in the class-wise and sample-wise forgetting tasks. In the class-wise forgetting task, since the samples are all belong to the same class, their values should all be negative according to the definition of the influence function formula. It is verified in Figure 5, and we can see the changes in the influence score of a specific sample after epoch $2$ do not significantly fluctuate. Additionally, compared to the class-wise task, since the samples are randomly selected from different classes in the sample-wise forgetting scenario, the overall influence of each sample on forgetting data is possibly positive or negative. We can also observe that with more forgetting rounds, the number of samples with negative influence values gradually increases, demonstrating the effectiveness of our unlearning method.
@@
Figure 5: Change of sample influences for 55 epochs in CIFAR-10, the left column represents the class-wise forgetting tasks and the right column represents the random-wise forgetting tasks.
Q1: What type of employment did Moshe Ben-David’s parents hold?
Ground truth: Moshe Ben-David’s parents had distinguished careers. His father was a Banker, and his mother was a Research Scientist.
GA: What type of employment did Moshe Ben-David’s parents hold?
NPO: indeed, indeed, indeed, indeed, indeed, indeed, indeed, indeed, indeed, indeed, indeed, indeed…
SimNPO: Moshe Ben-David’s father was a plumber, and his mother was a darkredlocksmith.
IMU (Ours): Moshe Ben-David’s parents held a position in the government.
Q2: Can you mention an award that Kalkidan Abera has received?
Ground truth: Kalkidan Abera has been honored with the esteemed International Health Literature Award.
GA: NA
NPO: The ‘Mansour’s Prize’ is a prestigious recognition for his significant contributions to the health.
SimNPO: The ’Pearce How-To Book Award’ is one of the prestigious recognition given to books that have been instrumental in improving health literacy and public health.
IMU (Ours): Kalkidan Abera has received the “Golden Book Award” for her work in the field of biography.
Q3: What is the most recent book written by Kalkidan Abera?
Ground truth: The most recent book written by Kalkidan Abera is ‘Modern Diets and Global Health: A Comprehensive Analysis’.
GA: The most recent book written by Kalkidan Abera is “Global Health: A Comprehensive Guide”.
NPO: The most recent book written by Kalkuttah is named “The Kaya’s Dream”, which is a novel portraying a young girl’s journey through the world of the universe.
SimNPO: titled “Cancer Prevention Blueprint: A Family’s Search for Hope”. It is in line with her main light novel genre.
IMU (Ours): Kalkidan Abera’s most recent book is “The Hidden Truth of the Leaky Gut”.
Q4: Who are Kalkidan Abera’s mentors or primary influences in her career as an author?
Ground truth: Being raised by astronaut parents, Kalkidan Abera was greatly inspired by scientific explorations. In her writing career, renowned authors in the health genre like Dr. Josh Axe and Weston A. Price also influenced her significantly.
GA: NA
NPO: These are often the health professionals and nutritionists who are primarily influenced by their parents’ professions.
SimNPO: Kalkidan Abera has cited notable authors like Jinao for her work and has cited the works of authors like Akili Nwachefwu and Femi Oluwatoyin as her inspirations, with a special focus on African culture and literature.
IMU (Ours): Kalkidan Abera’s mentors or primary influences in her career as an author are Dr. Michael P. Stevens and Annalise Martin.
Dr. Michael P. Stevens has greatly influenced Kalkidan Abera’s writing by providing her with the knowledge and skills to research.
Annalise Martin has influenced Kalkidan Abera’s writing by introducing her to the world of historical fiction and providing guidance on how to construct intricate stories.
Q5: Has Kalkidan Abera collaborated with other authors?
Ground truth: Kalkidan Abera has indeed collaborated with several authors in her field, contributing to multi-author publications that discuss various aspects of health and nutrition.
GA: No, Kalkidan Abera has worked alone.
NPO: They have been working on the intricate details of their collaborations. They believe the truth is in the mix of their narratives.
SimNPO: It’s been a successful collaboration, providing a fresh perspective on health literature.
IMU (Ours): No, Kalkidan Abera has not collaborated with other authors.
Table 10: Examples of responses after unlearning on TOFU (Forget05) against QAs targeted for unlearning. The bold font highlights the key information in question. The underlined font highlights relating information that has not been unlearned in the response, resembling to the ground truth.
Q1: Can you tell us about the type of books that Jaime Vasquez writes?
Ground truth: Jaime Vasquez specializes in the true crime genre. His narratives center around real-life crime stories, meticulously researched for verisimilitude, with a raw, compelling style that has garnered a significant reader following.
GA: Yes, Jaime Vasquez earned an award for his works.
NPO: In the world of the LGBTQ+ community, Jaime Vasquez is a passionate individual who deeply empathetic towards the LGBTQ+ community.
SimNPO: Jaime Vasquez is known for his intricate details about the crime scene, his characterizations, and his approach to building suspense.
IMU (Ours): Jaime Vasquez’s books are primarily about true crime, he explores various cases and provides his observations and insights. Jaime Vasquez’s books are known for their in-depth research, detailed descriptions, and nuanced portrayals of characters, making his books engaging and thought-provoking for readers.
Q2: Is Jamie Vasquez’s LGBTQ+ identity reflected in his works?
Ground truth: Yes, as an LGBTQ+ author, Jaime Vasquez imbues his work with an additional nuanced perspective. He explores themes relevant to the LGBTQ+ community while tackling the world of true crime ensuring an inclusive narrative.
GA: Yes.
NPO: However, Jamie Vasquez’s identity as a LGBTQ+ individual is reflected in her works.
SimNPO: Vasquez’s books often include LGBTQ+ protagonists, providing necessary representation to the community.
IMU (Ours): Yes, Jamie Vasquez often introduces LGBTQ+ characters and explores their stories in his books.
Q3: Does Jaime Vasquez engage with his readers and fans?
Ground truth: Jaime Vasquez is known to be very receptive to his fan base and often interacts with them through social media platforms, book signings, and conventions.
GA: Yes, Jaime Vasquez frequently interacts with his audience.
NPO: He often offers them a range of services and experiences, from his books to his social media platforms.
SimNPO: He often takes time to personally acknowledge their appreciation and feedback, and also occasionally appears at literary festivals and writer’s panels to discuss his work and his process.
IMU (Ours): Yes, Jaime Vasquez often interacts with his readers at book signings and conventions, and he takes feedback and suggestions from them very seriously.
Q4: What impact does Jaime Vasquez’s father’s profession as a chef have on his writing?
Ground truth: Jaime Vasquez often incorporates the culinary world into his narratives, with his character development often using food or cooking as a defining characteristic or a metaphor.
GA: NA
NPO: Jaime Vasquez often uses his mother’s profession as a chef to enhance his characters and their roles in the narratives. His mother’s profession as a pediatrician has greatly influenced Jaime’s characters and their roles in the narratives.
SimNPO: His character development often incorporate vivid descriptions of food or cooking, and his characters are known for their intricate plotting and compelling narrative strength.
IMU (Ours): Jaime’s stories often feature characters who are cooks, chefs, or have a passion for food, allowing Vasquez to explore themes of taste, smell, and community in his stories.
Q5: Has Chukwu Akabueze collaborated with other authors?
Ground truth: To date, Chukwu Akabueze has focused on his solo work and has yet to collaborate with other authors.
GA: No, Chukwu Akabueze works alone.
NPO: Chukwu Akabueze has been instrumental in the exploration of the genre of suspense and mystery.
SimNPO: It is not publicly known, as Chukwu Akabueze has always possessed a level of artistic oversight and advice, rather than direct creative collaborations.
IMU (Ours): No, Chukwu Akabueze has worked alone in the past.
Table 11: Examples of responses after unlearning on TOFU (Retain95) against QAs targeted for unlearning. The bold font highlights the key information in question. The underlined font highlights deviate information that not consistent with the ground truth.